OpenStudy (anonymous):

What is the equation for the axis of symmetry in the function? y = -2x2 + 5x - 4 y = 5/4 x = 5/4 y = -2/5 x = -2/5

jimthompson5910 (jim_thompson5910):

anything of the form y = ax^2 + bx + c has the axis of symmetry of x = -b/(2a)

OpenStudy (anonymous):

x=-5/-4

jimthompson5910 (jim_thompson5910):

reduce that

jimthompson5910 (jim_thompson5910):

what do you get when you do so?

OpenStudy (anonymous):

x=5/4

jimthompson5910 (jim_thompson5910):

good